OpenStudy (anonymous):

what is csc (2π/3)? explain

OpenStudy (anonymous):

Do you know what the Unit circle is?

OpenStudy (aravindg):

csc 2 pi/3= csx (pi-pi/3)

OpenStudy (aravindg):

*csc (pi-pi/3) =1/sin(pi-pi/3)

OpenStudy (aravindg):

can u do it now ?

OpenStudy (anonymous):

Go to the Unit circle and 2pi/3 = 120 degrees therefore you have to find out the value of csc at 120 degrees

OpenStudy (anonymous):

is it 1.15?

OpenStudy (anonymous):

http://web.gccaz.edu/~jgutier5/Tables/The%20Unit%20Circle%20Table%20Of%20Values.pdf If you look at this chart. you will find the value for csc at 2pi/3 which is csc 120 degrees. 2pi/3 is 120 because you take 2 and multiply it by 180 and divide by 3 to get 120

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

yes i see that part

OpenStudy (anonymous):

so csc of 120 degrees is 1.15470053838

OpenStudy (anonymous):

No you are looking for the exact value. you can't use your calculator to find it. you have to use the chart.

OpenStudy (anonymous):

are you looking for the exact value?

OpenStudy (anonymous):

i'm not sure what the question asks for

OpenStudy (anonymous):

Im pretty sure you are looking for the exact value.. therefore that link i gave you gives you the exact value.

OpenStudy (anonymous):

where is it on the chart?

OpenStudy (anonymous):

\[\frac{ 2\sqrt{3} }{ 3}\]

OpenStudy (aravindg):

1.15

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

look at 120 degrees and at the top it has csc cot sin cos tan all of those. And you will find it.

OpenStudy (aravindg):

you are right !

OpenStudy (anonymous):

it says 2sqrt3/3

OpenStudy (anonymous):

yes thats the answer.

OpenStudy (anonymous):

oh ok, thanks!!!

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

what about for sin(13π/4)? thats not on the chart is it?

OpenStudy (anonymous):

you have to find a coterminal angle for that. therefore you will be looking at 5pi/4 all i did was take 2pi from 13pi/4 \[\frac{ 13 }{ 4 }-\frac{ 2 }{ 1 }\]

OpenStudy (anonymous):

which gives you 5/4 which is 5pi/4.

OpenStudy (anonymous):

then 5x180 divided by 4 = 225 degrees so look for the exact value of 225 degrees for sine

OpenStudy (anonymous):

which should be \[\frac{ -\sqrt{2} }{ 2 }\]

OpenStudy (anonymous):

ooh thanks

OpenStudy (anonymous):

5π/4 13π/4-2π 13/4-2/1 5*180/4=225° -√(2)/2

OpenStudy (anonymous):

so is that correct or did i write it wrong

OpenStudy (anonymous):

you did it right. except what i did with the 13/4-2 was i found a coterminal angle for it. to do that you take away 2pi from 13/4 so 13/4-2 is just a regular subtraction of fractions.

OpenStudy (anonymous):

so it should be 13/4-2π?

OpenStudy (anonymous):

yes but what you did was correct... you might just want to show the teacher how you got 5pi/4 is what I'm saying.